Please help
Solve for x and y
4x-7y=8
3x-8y=5
I tried doing the elimination method by multiplying the top equation by -3 and the bottom by 4 so I could eliminate the x part but then I got stuck.

Answer by elima(1433)

4x-7y=8
3x-8y=5
(-3)4x-7y=8
-12x+21y=-24
===============
(4)3x-8y=5
12x-32y=20
=================
-12x+21y=-24
+12x-32y=20
-------------
-11y=-4
y=4/11
============================
Now plug that in as your y;
4x-7(4/11)=8
4x-28/11=8
4x=8+28/11
4x=88/11+28/11
4x=116/11
x=116/44=29/11
====================
Check;
4(29/11)-7(4/11)=8
116/11-28/11=8
88/11=8
8=8
:)